# A Survey on Wi-Fi Sensing Generalizability: Taxonomy, Techniques, Datasets, and Future Research Prospects

### v1

@@ 2. Taxonomy
Achieving generalizability in Wi-Fi sensing involves addressing challenges across different stages, including device deployment, signal processing, feature learning, and model deployment. While some studies focus on a single stage, others span multiple stages, tackling two or even three phases simultaneously. To provide a clear and systematic overview of existing research, we organize the surveyed works along the Wi-Fi sensing pipeline, covering every step from device setup to model deployment.
2.1. Experimental Setup
@@
At the experimental setup stage, some studies enhance the generalization capability of Wi-Fi sensing by adopting distributed antenna deployment, integrating additional devices, and collecting more extensive datasets.
2.1.1. Distributed Antennas
@@
WiNDR (Qin et al., 2024) and WiCross (Qin et al., 2023) position three antennas from the transmitter and three antennas from the receiver in a staggered, distributed manner around a 360-degree circle, with the subject performing gestures at the center. This configuration achieves direction-agnostic gesture recognition. Additionally, WiSDAR (Wang et al., 2018) tests various distributed antenna topologies, including line, hexagon, square, and random shapes, to explore regions that are most sensitive to human activities, achieving highly accurate and reliable recognition results.
2.1.2. Distributed Devices
@@
Some studies utilize additional WiFi devices to enhance sensing range and improve generalization to human orientation and position. For instance, ReWiS (Bahadori et al., 2022), Widar 3.0 (Zheng et al., 2019), and OneFi (Xiao et al., 2021) adopt configurations of 1 transmitter with 6 receivers and 1 transmitter with 4 receivers, respectively, to extract environment-independent features. WiTraj (Wu et al., 2021b) leverages 3 receivers placed at different viewing angles to capture human walking and achieve robust trajectory reconstruction. Additionally, Wang et al. (Wang et al., 2022a) propose a theoretical model analyzing how the distance between the transmitter and receiver influences the sensing range. Optimizing the placement of these devices can effectively enhance the sensing area and reduce the impact of environmental interference.
2.1.3. Scaling up dataset
@@
Scaling up training datasets is another effective approach to enhancing generalization capability, as demonstrated in Widar 3.0 (Zheng et al., 2019), Zhang et al. (Zhang et al., 2023h), MM-Fi(Yang et al., 2024a), and XRF55 (Wang et al., 2024b). For example, XRF55 (Wang et al., 2024b) shows that training on approximately 30,000 samples enables the sensing model to naturally achieve direction and position invariance, while also supporting adaptation to new environments through few-shot fine-tuning.
2.2. Signal Preprocessing Stage
@@
Many works focus on the signal preprocessing stage to extract domain-invariant features to enhance the generalizability of Wi-Fi sensing. These approaches can be broadly categorized into four types: (1) using signal analysis methods, such as Principal Component Analysis (PCA), to extract action-related features from raw signals; (2) processing phase or amplitude information to generate advanced action indicators; (3) extracting motion-specific indicators like Doppler Frequency Shift (DFS) and Body Velocity Profile (BVP) to capture dynamic movements; (4) employing antenna array techniques to derive measurements such as Angle of Arrival (AoA) and Angular Difference of Arrival (ADoA) for sensing. Next, we will introduce the papers corresponding to these four categories.
2.2.1. Signal Analysis Methods
@@
Signal analysis plays a crucial role in Wi-Fi sensing by extracting relevant features from raw signals, ensuring generalizability across various environments, locations, and orientations. WiHand (Lu et al., 2019) employs a Low Rank and Sparse Decomposition (LRSD) algorithm to separate gesture signals from background noise, enhancing its resilience to location variations. CLAR (Zhou et al., 2022a) utilizes Singular Spectrum Analysis (SSA) (Golyandina et al., 2001) to extract the trend components from Wi-Fi signals, enabling robust activity recognition. Gait-Enhance (Yang et al., 2023a) introduces a CSI reconstruction algorithm that mitigates the negative impact on recognition accuracy caused by varying CSI patterns associated with walking in different directions. Zhuo et al.  (Zhuo et al., 2023) apply a combination of Principal Component Analysis (PCA) and Variational Mode Decomposition (VMD) (Dragomiretskiy and Zosso, 2013) to reduce noise interference in non-ideal sleep positions, enabling position-free breath detection. Abuhoureyah et al. (Abuhoureyah et al., 2024) leverage Independent Component Analysis (ICA) and Continuous Wavelet Transform (CWT) to decompose CSI signals, improving multi-user human action recognition in complex environments.
2.2.2. Advanced Handcrafted Indicators
@@
Many researchers have focused on exploring robust handcrafted features for domain-independent sensing. WiHGR (Meng et al., 2021) uses a Phase Difference Matrix to extract phase-related features. PhaseAnti (Huang et al., 2021) introduces Nonlinear Phase Error Variation (NLPEV), which is independent of Cochannel Interference (CCI), to mitigate interference and enhance recognition accuracy. DPSense (Gao et al., 2022a) proposes the Error of Dynamic Phase Index (EDP-index) to evaluate the sensing quality of signal segments, prioritizing high-quality segments to improve gesture recognition performance. HandGest (Zhang et al., 2022d) combines Dynamic Phase Vector (DPV) and Motion Rotation Variable (MRV) for in-the-air handwriting recognition. Similarly, Yu et al. (Yu et al., 2023) extract dynamic phase features for position-independent gesture recognition. Additionally, AF-ACT (Zhang et al., 2022e) extracts semantic activity and temporal features to characterize activities across various locations. RoSeFi (Peng et al., 2023) identifies contextual association and waveform mirror symmetry in CSI data, introducing the Waveform Mirror Symmetry Factor (WMSF) to quantify symmetry between sit-stand postural transition (SPT) CSI data, aiding in sedentary behavior detection.
@@
Some works also extract features in the IQ complex plane. For instance, Wang et al. (Wang et al., 2024e) analyze the curvature of signal curves over time in the IQ plane. WiGesFree(Ding et al., 2024) introduces the Sample Point Distance (SPD), which uses changing CSI ratio patterns in the complex plane to capture gesture information accurately. LiKey (Peng et al., 2024) leverages rotational direction features in the CSI complex plane as an extraction of location-independent features. Other works transform Wi-Fi CSI data into images and use visual methods, such as pre-trained models and data augmentation, to improve generalization capabilities. Bu et al. (Bu et al., 2018) convert the amplitude of each CSI stream into a grayscale CSI image and fine-tune pre-trained networks like VGG16 and VGG19 for feature extraction. Zhang et al. (Zhang et al., 2023e) convert one-dimensional CSI time series into two-dimensional recurrence plots (RP) and apply data augmentation techniques like horizontal flipping, cropping, and color distortion to improve model performance and robustness in action recognition.
2.2.3. Motion Indicators
@@
Many works compute domain-independent motion indicators in speed, velocity, direction, and trajectory, to enhance the generalizability of Wi-Fi sensing systems. Sharp (Meneghello et al., 2022), OneSense (Zhao et al., 2024d), Yin et al. (Yin et al., 2021), Wi-learner (Feng et al., 2022), Niu et al. (Niu et al., 2021), and LAGER (Chen et al., 2024a) focus on utilizing Doppler Frequency Shift (DFS). Mini-Batch Alignment (van Berlo et al., 2023, 2022) incorporates both DFS and Gramian Angular Difference Field (GADF). WiLife (Li et al., [n. d.]) further explores Doppler velocity as a key motion indicator. WiTraj (Wu et al., 2021b) identifies speed ambiguity in the Doppler-MUSIC algorithm’s spectrum. To address this, it employs multiple receivers at different viewing angles and leverages the CSI quotient to eliminate this ambiguity, enabling robust trajectory reconstruction.
@@
Widar 3.0 (Zheng et al., 2019) introduces the Body-coordinate Velocity Profile (BVP), which captures the power distribution across different velocities of body parts involved in gesture movements. This approach enhances generalization to new locations, environments, and subjects while maintaining performance consistency. BVP and its variants have demonstrated remarkable generalization capabilities. For instance, Zhang et al. (Zhang et al., 2023i) and WiTransformer (Yang et al., 2023c) adopt BVP, while Shi et al. (Shi et al., 2023) and Bulugu (Bulugu, 2023) extend it to 3D BVP. GaitID (Zhang et al., 2020d) utilizes gait-specific BVP (gait-BVP), OneFi (Xiao et al., 2021) employs velocity distribution, and XFall (Chi et al., 2024) incorporates a speed distribution profile to enhance robustness. In addition to BVP, WiGesture (Gao et al., 2021) proposes the Motion Navigation Primitive (MNP) to improve gesture recognition accuracy. Shi et al. (Shi et al., 2020a) convert CSI amplitude and relative amplitude into frequency-domain representations, such as STFT holograms, to characterize the moving speeds of different body parts. Wang et al. (Wang et al., 2022c) use phase changes of the reflection path to relate the speed vector and gesture trajectory in the body coordinate system, implicitly capturing the impact of position and orientation, which is estimated by a preamble gesture in each new environment.
@@
The Fresnel Zone Model in physics is also widely used to enhance the generalization capability of Wi-Fi sensing. Widir (Wu et al., 2016) utilizes it to capture walking direction. Wang et al.(Wang et al., 2016e) apply the model to detect breast motion for respiration detection. WiNDR(Qin et al., 2024) leverages the model for direction-agnostic gesture recognition. Additionally, WiDIGR (Zhang et al., 2019a) and Wi-PIGR (Zhang et al., 2021d) use the Fresnel Zone Model to eliminate direction information for gait recognition. Statistical EM methods have also been applied to Wi-Fi sensing. WiDetect (Zhang et al., 2019b) uses statistical EM theory to model motion detection, connecting the Autocorrelation Function (ACF) of CSI to target motion, independent of environment, location, orientation, and subjects. Zhu et al. (Zhu et al., 2024) extract the Amplified ACF spectrum from CSI, enabling motion features to be isolated, independent of environmental factors.
2.2.4. Angle Indicators
@@
Angle indicators are also used to enhance the generalizability of Wi-Fi sensing. Li et al. (Li et al., 2020) use Angle Difference of Arrival (ADoA) for location-free CSI-based activity recognition. EasyTrack (Han et al., 2023) utilizes AoA–AoD values, leveraging changes in the length and direction of the Wi-Fi signal path. Through incremental calculations and dynamic path extraction, it can reconstruct gesture trajectories in real-time without relying on the initial gesture position or a fixed device location. 3D-ID (Ren and Yang, 2022) uses 2D AoA estimation of signal reflections to enable Wi-Fi to visualize a person in the environment. This visualization is then digitized into a 3D body representation using deep learning, which extracts both the static body shape and dynamic walking patterns for person re-identification.
2.3. Feature Learning Stage

### v2

@@ II Taxonomy
To enhance the generalization of Wi-Fi sensing systems, researchers must address challenges across various stages, including device deployment, signal processing, feature learning, and model deployment. While some studies concentrate on a single stage, others tackle multiple stages simultaneously, addressing two or three phases at once. To provide a clear and systematic overview of existing research, we categorize the surveyed works based on the Wi-Fi sensing pipeline, as illustrated in Fig. 4.
II-A Experimental Setup
@@
At the experimental setup stage, some studies enhance the generalization capability of Wi-Fi sensing by adopting distributed antenna deployment, integrating additional devices, and collecting more extensive datasets, as shown in Fig. 5.
II-A1 Distributed Antennas
@@
WiNDR [63] and WiCross [64] position three antennas from the transmitter and three antennas from the receiver in a staggered, distributed manner around a 360-degree circle, with the subject performing gestures at the center. This configuration achieves direction-agnostic gesture recognition. Additionally, WiSDAR [65] tests various distributed antenna topologies, including line, hexagon, square, and random shapes, to explore regions that are most sensitive to human activities, achieving highly accurate and reliable recognition results.
II-A2 Distributed Devices
@@
Some studies utilize additional WiFi devices to enhance sensing range and improve generalization to human orientation and position. For instance, ReWiS [66], Widar 3.0 [67], and OneFi [68] adopt configurations of 1 transmitter with 6 receivers and 1 transmitter with 4 receivers, respectively, to extract environment-independent features. WiTraj [69] leverages 3 receivers placed at different viewing angles to capture human walking and achieve robust trajectory reconstruction. Additionaly, Wang et al. [245] propose a theoretical model analyzing how the distance between the transmitter and receiver influences the sensing range. Optimizing the placement of these devices effectively enhances the sensing area and reduces the impact of environmental interference.
II-A3 Scaling up dataset
@@
Scaling up training datasets is another effective approach to enhancing generalization capability, as demonstrated in Widar 3.0 [67], Zhang et al. [71], MM-Fi [72], XRF55 [70] and CSI-Bench [73]. For example, XRF55 [70] shows that training on approximately 30,000 samples enables the sensing model to naturally achieve direction and position invariance, while also supporting adaptation to new environments through few-shot fine-tuning.
Figure 6: Signal Preprocessing Stage
II-B Signal Preprocessing Stage
@@
Many works focus on the signal preprocessing stage to extract domain-invariant features to enhance the generalizability of Wi-Fi sensing. These approaches can be broadly categorized into four types: (1) using signal analysis methods, such as Principal Component Analysis (PCA), to extract action-related features from raw signals; (2) processing phase or amplitude information to generate advanced action indicators; (3) extracting motion-specific indicators like Doppler Frequency Shift (DFS) and Body Velocity Profile (BVP) to capture dynamic movements; (4) employing antenna array techniques to derive measurements such as Angle of Arrival (AoA) and Angular Difference of Arrival (ADoA) for sensing. In the following, we introduce four categories respectively.
II-B1 Signal Analysis Methods
@@
Signal analysis plays a crucial role in Wi-Fi sensing by extracting relevant features from raw signals, ensuring generalizability across various environments, locations, and orientations. WiHand [74] employs a Low Rank and Sparse Decomposition (LRSD) algorithm to separate gesture signals from background noise, enhancing its resilience to location variations. CLAR [75] utilizes Singular Spectrum Analysis (SSA) [246] to extract the trend components from Wi-Fi signals, enabling robust activity recognition. Gait-Enhance [76] introduces a CSI reconstruction algorithm that mitigates the negative impact on recognition accuracy caused by varying CSI patterns associated with walking in different directions. Zhuo et al.  [77] apply a combination of Principal Component Analysis (PCA) and Variational Mode Decomposition (VMD) [247] to reduce noise interference in non-ideal sleep positions, enabling position-free breath detection. Abuhoureyah et al. [78] leverage Independent Component Analysis (ICA) and Continuous Wavelet Transform (CWT) to decompose CSI signals, improving multi-user human action recognition in complex environments.
II-B2 Advanced Handcrafted Indicators
@@
Many works have focused on exploring robust handcrafted features for domain-independent sensing. WiHGR [79] uses a Phase Difference Matrix to extract phase-related features. PhaseAnti [80] introduces Nonlinear Phase Error Variation (NLPEV), which is independent of Cochannel Interference (CCI), to mitigate interference and enhance recognition accuracy. DPSense [81] proposes the Error of Dynamic Phase Index (EDP-index) to evaluate the sensing quality of signal segments, prioritizing high-quality segments to improve gesture recognition performance. HandGest [82] combines Dynamic Phase Vector (DPV) and Motion Rotation Variable (MRV) for in-the-air handwriting recognition. Similarly, Yu et al. [83] extract dynamic phase features for position-independent gesture recognition. Additionally, AF-ACT [84] extracts semantic activity and temporal features to characterize activities across various locations. RoSeFi [85] identifies contextual association and waveform mirror symmetry in CSI data, introducing the Waveform Mirror Symmetry Factor (WMSF) to quantify symmetry between sit-stand postural transition (SPT) CSI data, aiding in sedentary behavior detection.
@@
Some works also extract features in the IQ complex plane. For instance, Wang et al. [86] analyze the curvature of signal curves over time in the IQ plane. WiGesFree[87] introduces the Sample Point Distance (SPD), which uses changing CSI ratio patterns in the complex plane to capture gesture information accurately. LiKey [88] leverages rotational direction features in the CSI complex plane as an extraction of location-independent features. Other works transform Wi-Fi CSI data into images and use visual methods, such as pre-trained models and data augmentation, to improve generalization capabilities. Bu et al. [89] convert the amplitude of each CSI stream into a grayscale CSI image and fine-tune pre-trained networks like VGG16 and VGG19 for feature extraction. Zhang et al. [90] convert one-dimensional CSI time series into two-dimensional recurrence plots (RP) and apply data augmentation techniques like horizontal flipping, cropping, and color distortion to improve model performance and robustness in action recognition.
II-B3 Motion Indicators
@@
Many works compute domain-independent motion indicators in speed, velocity, direction, and trajectory, to enhance the generalizability of Wi-Fi sensing systems. Sharp [91], OneSense [92], Yin et al. [93], Wi-learner [94], Niu et al. [95], and LAGER [96] focus on utilizing Doppler Frequency Shift (DFS). Mini-Batch Alignment [98, 99] incorporates both DFS and Gramian Angular Difference Field (GADF). WiLife [97] further explores Doppler velocity as a key motion indicator. WiTraj [69] identifies speed ambiguity in the Doppler-MUSIC algorithm’s spectrum. To address this, it employs multiple receivers at different viewing angles and leverages the CSI quotient to eliminate this ambiguity, enabling robust trajectory reconstruction.
@@
Widar 3.0 [67] introduces the Body-coordinate Velocity Profile (BVP), which captures the power distribution across different velocities of body parts involved in gesture movements. This approach enhances generalization to new locations, environments, and subjects while maintaining performance consistency. BVP and its variants have demonstrated remarkable generalization capabilities. For instance, Zhang et al. [100] and WiTransformer [101] adopt BVP, while Shi et al. [102] and Bulugu [103] extend it to 3D BVP. GaitID [30] utilizes gait-specific BVP (gait-BVP), OneFi [68] employs velocity distribution, and XFall [104] incorporates a speed distribution profile to enhance robustness. In addition to BVP, WiGesture [105] proposes the Motion Navigation Primitive (MNP) to improve gesture recognition accuracy. Shi et al. [106] convert CSI amplitude and relative amplitude into frequency-domain representations, such as STFT holograms, to characterize the moving speeds of different body parts. Wang et al. [107] use phase changes of the reflection path to relate the speed vector and gesture trajectory in the body coordinate system, implicitly capturing the impact of position and orientation, which is estimated by a preamble gesture in each new environment.
@@
The Fresnel Zone Model describes the signal propagation process with better interpretability, and is therefore widely used to enhance the generalization capability of Wi-Fi sensing. Widir [108] utilizes it to capture walking direction. Wang et al. [109] apply the model to detect breast motion for respiration detection. WiNDR [63] leverages the model for direction-agnostic gesture recognition. Additionally, WiDIGR [110] and Wi-PIGR [111] use the Fresnel Zone Model to eliminate direction information for gait recognition. Statistical EM methods have also been applied to Wi-Fi sensing. WiDetect [112] uses statistical EM theory to model motion detection, connecting the Autocorrelation Function (ACF) of CSI to target motion, independent of environment, location, orientation, and subjects. Zhu et al. [113] extract the Amplified ACF spectrum from CSI, enabling motion features to be isolated, independent of environmental factors.
II-B4 Angle Indicators
@@
Angle indicators are also used to enhance the generalizability of Wi-Fi sensing. Li et al. [114] use Angle Difference of Arrival (ADoA) for location-free CSI-based activity recognition. EasyTrack [115] utilizes AoA–AoD values, leveraging changes in the length and direction of the Wi-Fi signal path. Through incremental calculations and dynamic path extraction, it can reconstruct gesture trajectories in real-time without relying on the initial gesture position or a fixed device location. 3D-ID [116] uses 2D AoA estimation of signal reflections to enable Wi-Fi to visualize a person in the environment. This visualization is then digitized into a 3D body representation using deep learning, which extracts both the static body shape and dynamic walking patterns for person re-identification.
II-C Feature Learning Stage